OpenStudy (anonymous):

x + 9y=4 (1) x=3-9y (2)

jhonyy9 (jhonyy9):

what is the question ?

OpenStudy (anonymous):

Solve the folowing systems of equations. x + 9y=4 (1) x=3-9y (2) What is the solution of the system?

jhonyy9 (jhonyy9):

so than there is x+9y=4 and x=3-9y - from second equation can you make the same organized by terms like first equation please ?

jhonyy9 (jhonyy9):

this is the first step

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

wait

OpenStudy (anonymous):

ok